my younger sister's age today is 3 times , what it will be 3 yrs from now minus 3 times what her age was 3 yrs ago. find her present age

Grade:7

1 Answers

Saswat Pani
6 years ago
Let her present age is x 3 yrs from now minus 3 times what her age was 3 yrs ago = 3(x+3) -3(x-3)= x3x - 3x +9 + 9 = xX= 18